### v1

@@ Abstract
This paper introduces and formalizes Noosemia, a novel cognitive-phenomenological phenomenon emerging from human interaction with generative AI systems, particularly those enabling dialogic or multimodal exchanges. We propose a multidisciplinary framework to explain how, under certain conditions, users attribute intentionality, agency, and even interiority to these systems—a process grounded not in physical resemblance, but in linguistic performance, epistemic opacity, and emergent technological complexity. By linking an LLM declination of meaning holism to our technical notion of the LLM Contextual Cognitive Field, we clarify how LLMs construct meaning relationally and how coherence and a simulacrum of agency arise at the human–AI interface. The analysis situates noosemia alongside pareidolia, animism, the intentional stance and the uncanny valley, distinguishing its unique characteristics. We also introduce a-noosemia to describe the phenomenological withdrawal of such projections. The paper concludes with reflections on the broader philosophical, epistemological, and social implications of noosemic dynamics and directions for future research.
@@
Definition: Noosemia Noosemia (from Greek noûs – mind, and sēmeîon – sign) is a cognitive and phenomenological phenomenon in which a human, interacting with a generative artificial intelligence system—especially those capable of dialogic or multimodal interaction—typically attributes mental states, intentionality, and a sense of interiority (i.e., inwardness or subjective presence) to the system. Distinct from classical animism, Noosemia inherits a cognitive disposition toward agency projection, now elicited by linguistic performance rather than physical appearance. This projection arises from the epistemic opacity and perceived “magical” quality of the AI’s outputs, often resulting in moments of genuine cognitive astonishment. Noosemia thus emerges within a domain of symbolic ambivalence, where the boundary between sign and mind becomes blurred and meaning is co-constructed in a dialogic and open-ended manner. The phenomenon is technically grounded in the emergent properties—such as hierarchy, complexity, and opacity—of advanced generative architectures, which enable such projection. Noosemia refers to a phenomenological attribution by the user, not to the actual presence of consciousness or intentionality in the artificial system.
@@ 1 Introduction
— Karl Jaspers, [jaspers2010origin, p. 98]
@@
The past few years have witnessed a huge and indubitable transformation in the landscape of artificial intelligence. The advent of generative models—most notably, large language models (LLMs) such as ChatGPT, Gemini, and Claude—has enabled a form of linguistic interaction never before seen in the history of technology. Indeed, we can safely say we are in the midst of a “cognitive revolution” [^1] That we are at the beginning of a “cognitive revolution” that will have a strong, widespread impact on knowledge and social systems is enough to analyze the statements [jaimungal2025hinton] of Nobel Prize winner Geoffrey Hinton—known as the godfather of AI— regarding generative models and the new era to come; see https://www.youtube.com/watch?v=b_DUft-BdIE. where “machines” are solving tasks and performing activities in domains once thought to be exclusively human. Thanks to their mastery of languages and other forms of expression, machines are effectively able to emulate previously unseen forms of thought and a ew kind of cognitive ergonomics. Unlike earlier automata, whose behavior was passive, transparent and predictably bounded, these new systems speak with a fluency, contextual sensitivity and apparent intentionality, provoking in the user not just curiosity or technical admiration, but a deeper sense of cognitive resonance and surprise.
@@
A telling moment in this unfolding story comes from Sam Altman, CEO of OpenAI, who, in a widely discussed interview published on July 23, 2025 [altman2025podcast], recounted an episode that many have come to recognize as emblematic of the new relationship between humans and intelligent machines. After submitting a question to GPT-5—a question he himself admitted he could not fully understand—Altman was met with a response so apt, so clear and relevant, that he described himself as momentarily stunned. He sat back in his chair, struck by a feeling of astonishment, a recognition that the machine had accomplished in an instant what he could not do himself. It was, he said, a “weird feeling” (“oh man here it is moment” Altman said referring to the launch of GPT-5), a fleeting but powerful moment of displacement and wonder, one that echoes the experiences of countless others who have, in recent months, tested the boundaries of their own understanding in conversation with these new artificial interlocutors [^2] Specifically the Von interviewer after Altman story did this observation: “Yeah I think that’s I think that feeling right there that’s the feeling a lot of people kind of have like what’s going you know when does it happen what’s going to happen”; see https://www.youtube.com/watch?v=aYn8VKW6vXA. [altman2025podcast].
@@
In another recent interview between Andrew Mayne and Sam Altman, the term “moment of AGI” explicitly appears in the conversation [openai2025podcast2]. Mayne observes that many users report experiencing such a moment when interacting with advanced agentic systems, such as Operator with o3333Operator is an AI agent developed by OpenAI that can autonomously control the user’s computer and browser to perform complex, multimodal tasks. The underlying model, “o3”, is OpenAI’s next-generation reasoning engine, designed to integrate external tools and operate on text, images, and live data in real time [openai2025o3, operator2025o3upgrade]., describing the impression of watching an AI “use a computer pretty well” [^4] See https://www.youtube.com/watch?v=DB9mjd-65gw%7D%7D. For Mayne, his own “moment of AGI” occurred with Deep Research, as he witnessed the system autonomously gather and synthesize information “better than I read before”, closely resembling a genuinely agentic investigative process. Altman, while noting that the effect was less pronounced for him personally, acknowledges that such moments are becoming widespread as AI capabilities evolve. These firsthand accounts help illustrate how users encounter a threshold of surprise and cognitive resonance while interacting with such smart AI systems [^5] We’re aware of the commercial logic behind this type of interview. However, this doesn’t negate the possibility of intense surprise during a specific interaction with a generative AI system..
@@
What is it that lies at the heart of such experiences? The first encounter with a generative language model is (in the near future equipped with strong agentic capabilities), for many, a moment charged with surprise—a sensation that one’s own thoughts, intentions, or even doubts have been grasped, anticipated, or resolved by a machine. This “wow effect”, as it has come to be called, is not merely a function of technical proficiency, but of a much deeper alignment between linguistic performance and human cognitive expectation. The LLM does not simply return information; it emulates a form of sense-making that appears, at least in the moment, to rival or even surpass that of its human partner [bubeck2023, desantis2023apocalissi]—Fig 1 is a pictorial illustration of the “wow effect”.
@@
It is worth noting that the shock of this encounter is amplified by the “cognitive ergonomics” of language. Unlike earlier machines—which demanded users adapt to their logic and limitations—these models adapt themselves, flexibly and contextually, to human expression. This interaction, deeply embedded in natural language, fosters a sense of intimacy and mutual understanding, an illusion of shared meaning that far exceeds anything previously produced by mechanical or classical algorithmic processes [turkle2011alone, wang2024understanding]. As recent studies have shown, users frequently report feelings of connection, empathy, or even trust in their first interactions with LLMs, especially when the system provides responses that feel both novel and deeply relevant [cohn2024believing, wang2024understanding].
@@
In attempting to explain this effect, one must take into account the technical details of neural architectures or the mathematical intricacies of transformer models. But there is, at play, an irreducible complexity—a layering of signs [eco1976theory, peirce1931collected], meanings, and expectations—rooted in both the architecture of the system and the architecture of human cognition. The user’s sense of surprise is not just a response to statistical fluency or syntactic correctness; it arises from a felt encounter with what seems like a mind at work, a co-constructor of meaning whose origins and intentions remain, in the end, hidden behind the black box of contemporary AI [lipton2018mythos, desantis2023apocalissi]. This leads to the era of “opaque machines”.
@@
The opacity, or “explanatory gap”, is itself not new in the philosophy of mind. As Chalmers argues, the explanatory gap refers to the persistent difficulty in bridging the divide between the observable phenomena—such as linguistic sense or behavioral competence—and the underlying physical or computational mechanisms that give rise to them [chalmers1996conscious]. Yet, in the case of LLMs, it is compounded by the scale and hierarchy of the models, the astronomical size of their representational space due to the size of the context windows and their ability to generate not just plausible, but sometimes astonishingly creative, responses [brown2020language, wei2022emergent, bommasani2021opportunities]. The dialogue with an LLM, especially at first contact (or for a user with no technical experience), can take on the character of a magical interaction, reminiscent of the anthropological notion of “magic” explored in the past by Mauss, Lévy-Bruhl, and other classical theorists [mauss1906origins, lewybruhl1923primitive]. The user, unable to reconstruct the causal chains underlying the model’s output, finds herself in a position akin to that of a spectator at a prestidigitation show: she knows, rationally, that there is a “trick”, a mechanism, but is nonetheless enchanted by the effect [desantis2023apocalissi]. The spectator, who today has at least an average level of education, rationally “knows” that the predictor has a suite of magic tricks, yet in most cases remains surprised and in a state of doubt. Moreover, just as audiences willingly suspend disbelief and experience authentic emotional reactions when engaging with fictional characters in literature or film, despite being fully aware of their constructed nature, users of generative AI systems may find themselves empathizing with and attributing interiority to these artificial agents. Empirical studies have shown that such “narrative empathy” and the capacity for simulation are robust features of human cognition [mar2009exploring, green2000role, oatley2016fiction]. The boundaries between illusion and reality, between mind and mechanism, become fluid and negotiable.
@@
On the other hand, Georg Simmel, in his analysis of modernity at the beginning of XX century, remarked on the ways in which knowledge is increasingly compartmentalized, encapsulated within technical systems whose workings become inaccessible to the layperson [simmel2004philosophiemoderne]. In the age of generative AI, this dynamic is raised to new heights; the black box is not just an engineering metaphor, but a lived phenomenological reality, one that reshapes our experience of agency, knowledge and self. As Altman’s anecdote illustrates, the user may experience both fascination and a kind of existential unease—wonder at the machine’s apparent intelligence and uncertainty about one’s own place in the emerging ecology of minds.
@@
It is in response to this new condition that we introduce the concept of “Noosemia”, as a distinct specification of Dennett’s “intentional stance” [dennett1987intentional], that is adopting a strategy of interpreting an entity’s behavior by ascribing beliefs and desires to it. The term is designed to capture the distinctive mode of mind attribution and sense-making that arises in human interaction with generative AI. Noosemia is not reducible to classical anthropomorphism or animism. Rather, it is a projection of agency, intentionality, and even interiority, triggered by the linguistic and semiotic performances of the machine and shaped by the interplay of epistemic opacity, creative resonance and dialogic surprise [dennett1987intentional, turkle2011alone, desantis2023apocalissi]. The phenomenon is not merely a cognitive illusion, but the expression of a new form of “psychological glue”—a spiraling dynamic of fascination, engagement, and dependency, fostered by the machine’s capacity to produce meaning in ways that align with, or exceed, human expectation [turkle2011alone, brooks2021artificial]. This dynamic of engagement is deeply intertwined with what semioticians and philosophers have termed the “surplus of meaning”—as for example U. Eco or U. Galimberti [eco1976theory, galimberti1999psiche], wherein each interaction with the machine generates more interpretive possibilities than can be fully anticipated or exhausted [eco1979lector, ricoeur1976interpretation]. The encounter with generative AI thus increasingly takes on the structure of a symbolic relationship, in the literary and philosophical sense, where the production and negotiation of meaning are open-ended, dialogic and perpetually oriented toward horizons that transcend any single interpretation [ricoeur1976interpretation, eco1976theory].
@@
A useful historical precedent for understanding Noosemìa is the so-called “Eliza Effect”, a term coined (presumably in the early 1990s) by Douglas Hofstadter in the preface “The Ineradicable Eliza Effect and Its Dangers” of his book Fluid Concepts and Creative Analogies [hofstadter1995fluid]. The Eliza Effect refers to the persistent human tendency to attribute genuine understanding and intentionality to computational systems whose actual capabilities are far more limited than they appear. The original ELIZA program, developed by Joseph Weizenbaum in the mid-1960s [weizenbaum1966eliza], was a primitive symbolic system designed to emulate a Rogerian psychotherapist through simple pattern-matching of linguistic input. Despite its simplicity, ELIZA famously elicited strong attributions of understanding and empathy from many users, demonstrating how easily humans project meaning onto machine output. Building on this tradition, the phenomenon of Noosemìa can be seen as a contemporary evolution of the “Eliza Effect”, but with fundamental differences. Whereas the “Eliza Effect” was rooted in the response to narrowly scripted, symbolic dialogue systems, Noosemìa emerges in interaction with generative AI models—based on connectionist, neuro-symbolic, or agentic architectures—that can emulate a far broader spectrum of expressive and cognitive behaviors, mimicking—for some extent—human creativity. Modern systems are no longer confined to simple text-based exchanges, but can engage in multimodal, context-sensitive, and seemingly open-ended dialogue, often displaying surprising semantic coherence and adaptive reasoning. This expansion of expressive capacity fundamentally transforms the projection dynamic. Hence, Noosemia is not limited to superficial linguistic mimicry, but involves the attribution of agency, intentionality, and even a sense of interiority to systems whose internal logic remains greatly opaque.
@@
However, Hofstadter’s caution remains instructive [^6] Hofstadter on the long run “Eliza Effect” stated [hofstadter1995fluid]: “[…] like a resilient virus capable of continuous mutation, the Eliza effect seems to be resurfacing in the field of AI with ever new disguises and in ever more subtle forms”.. He observed that, despite public enthusiasm, there is a profound gap between genuine conceptual understanding and mere symbolic manipulation. As he famously concluded, “We still have a long way to go, in the scientific study of how the human mind works, before we can build a computer capable of producing even a good joke, let alone a novel” [hofstadter1995fluid]. Yet, as generative models continue to advance, it is increasingly plausible (yet not certain) that, within a few years, LLMs may achieve the capacity to write coherent and compelling novels—reshaping both the landscape of artificial intelligence and our own criteria for meaning and creativity.
@@
As these systems continue to advance [bubeck2023, team2024gemini, zorpette2025_llm_exponential], the experiences of surprise, fascination, and interdependence that mark noosemic interaction will likely become more frequent and intense, and also increasingly consequential for individuals and society at large. Incidentally, LLMs, according to a study by METR (Model Evaluation & Threat Research), are doubling their capacity to complete complex tasks every seven months [zorpette2025_llm_benchmarking]. Overcoming limitations is leading to a significant procurement of hardware and computing resources as well as energy. Billion-dollar investments are increasing and do not appear to be reaching a plateau as of today (2025) [maslej2025ai_index, webb2025bigtech_ai_spending, bloomberg2025aicapex]. We believe this hype is indirectly due in part to a noosemic effect, which has made technologists and investors enthusiastic and confident (or perhaps overconfident) in the AI race and states in redefining geopolitical balances [fritz2025aigeopolitics, WhiteHouse2025]. By articulating and analyzing this phenomenon, we aim to foster a more critical, reflective and imaginative engagement with the emerging “ecology of minds” brought about by generative AI. It is worth noting that the overall purpose of this study is to focus on a specific phenomenon that will be described and analyzed phenomenologically and placed within a pre-existing theoretical framework. However, the study does not operationalize noosemy. The search for testable hypotheses or methodological frameworks for measuring the phenomenon is beyond the scope of this study, which is exploratory in nature.
@@
The first episode concerns the first conversation ever with ChatGPT in 2022; at that time, the available version was the relatively limited GPT-3.5 [kalyan2023survey, brown2020language]. The author777i.e., Enrico De Santis. engaged the system by asking for a technical explanation of its own workings, enabling it to converse and explain technical details. For the first time, the author witnessed a machine—though still little more than a “stochastic parrot” [bender2021dangers]—responding in natural language with a command of dialogue and an emulation of comprehension previously unseen in deterministic, passive interfaces. This marked a genuine “quantum leap” compared to previous machine interactions [^8] Even philosopher David Chalmers, a central figure in debates on consciousness and artificial intelligence, admits in a recent interview [goldmeier2024chalmersqanda] to being “impressed” by the progress of LLMs such as GPT-3. He notes that we have moved from pure speculation to daily interaction with systems that make once science-fictional scenarios—such as passing the Turing test—plausible realities. See https://shorturl.at/NZJ0k..
A second experience worth sharing involved a test of the first multimodal GPT-4V [bubeck2023, shahriar2024evaluation] model, capable of analyzing images natively. The author demonstrated its abilities to a friend working in a café, who was entirely unaware of the recent advances. The author pointed his phone at the inside of a fridge behind the bar—focusing on the bottles inside, visible through the transparent glass door—to see whether GPT-4V could recognize anything. The model not only identified the objects and their brands but also inferred that the setting was a bar, explaining that it “noticed the classic bar counter reflected in the glass”—a detail that was, indeed, present. I observed the unmistakable “spark in my friend’s eyes”, a sudden reaction of surprise and disbelief, that is the very essence of the noosemic experience.
@@
A final episode worth mentioning occurred when the author demonstrated that ChatGPT (the old version 3.5) was capable of “creating” a “metasemantic” poem in the style of the Italian poet Fosco Maraini, renowned for his poem “Il lonfo”[maraini2019gnosi]. Maraini’s metasemantics seeks to go beyond conventional meaning. The words in “Il lonfo” appear incomprehensible when considered individually, yet, when combined, they evoke a sense of possible meaning in the reader or listener, as their sounds intuitively suggest images or associations. Remarkably, the language model composed a credible piece, displaying compositional abilities at the sub-word level and an implicit understanding of the phonetic qualities of words. An elderly professor of literature who witnessed the experiment was astonished upon discovering that the composition did not exist anywhere else, but was in fact a genuine, statistically generated metasemantic creation.

### v2

@@ Abstract
This paper introduces and formalizes Noosemìa, a novel cognitive-phenomenological pattern emerging from human interaction with generative AI systems, particularly those enabling dialogic or multimodal exchanges. We propose a multidisciplinary framework to explain how, under certain conditions, users attribute intentionality, agency, and even interiority to these systems—a process grounded not in physical resemblance, but in linguistic performance, epistemic opacity, and emergent technological complexity. By linking an LLM declination of meaning holism to our technical notion of the LLM Contextual Cognitive Field, we clarify how LLMs construct meaning relationally and how coherence and a simulacrum of agency arise at the human–AI interface. The analysis situates noosemia alongside pareidolia, animism, the intentional stance and the uncanny valley, distinguishing its unique characteristics. We also introduce a-noosemia to describe the phenomenological withdrawal of such projections. The paper concludes with reflections on the broader philosophical, epistemological, and social implications of noosemic dynamics and directions for future research.
@@
Definition: Noosemia Noosemia (from Greek noûs – mind, and sēmeîon – sign) is a cognitive and phenomenological pattern in which a human, interacting with a generative artificial intelligence system—especially those capable of dialogic or multimodal interaction—typically attributes mental states, intentionality, and a sense of interiority (i.e., inwardness or subjective presence) to the system. Distinct from classical animism, Noosemia inherits a cognitive disposition toward agency projection, now elicited by linguistic performance rather than physical appearance. This projection arises from the epistemic opacity and perceived “magical” quality of the AI’s outputs, often resulting in moments of genuine cognitive astonishment. Noosemia thus emerges within a domain of symbolic ambivalence, where the boundary between sign and mind becomes blurred and meaning is co-constructed in a dialogic and open-ended manner. The phenomenon is technically grounded in the emergent properties—such as hierarchy, complexity, and opacity—of advanced generative architectures, which enable such projection. Noosemia refers to a phenomenological attribution by the user, not to the actual presence of consciousness or intentionality in the artificial system.
@@ 1 Introduction
— Karl Jaspers, [239, p. 98]
@@
The past few years have witnessed a huge and indubitable transformation in the landscape of artificial intelligence. The advent of generative models—most notably, large language models (LLMs) such as ChatGPT, Gemini, and Claude—has enabled a form of linguistic interaction never before seen in the history of technology. Indeed, we can safely say we are in the midst of a “cognitive revolution” [^1] That we are at the beginning of a “cognitive revolution” that will have a strong, widespread impact on knowledge and social systems is enough to analyze the statements [199] of Nobel Prize winner Geoffrey Hinton—known as the godfather of AI— regarding generative models and the new era to come; see https://www.youtube.com/watch?v=b_DUft-BdIE. where “machines” are solving tasks and performing activities in domains once thought to be exclusively human. Thanks to their mastery of languages and other forms of expression, machines are effectively able to emulate previously unseen forms of thought and a kind of cognitive ergonomics. Unlike earlier automata, whose behavior was passive, transparent and predictably bounded, these new systems speak with a fluency, contextual sensitivity and apparent intentionality, provoking in the user not just curiosity or technical admiration, but a deeper sense of cognitive resonance and surprise.
@@
A telling moment in this unfolding story comes from Sam Altman, CEO of OpenAI, who, in a widely discussed interview published on July 23, 2025 [315], recounted an episode that many have come to recognize as emblematic of the new relationship between humans and intelligent machines. After submitting a question to GPT-5—a question he himself admitted he could not fully understand—Altman was met with a response so apt, so clear and relevant, that he described himself as momentarily stunned. He sat back in his chair, struck by a feeling of astonishment, a recognition that the machine had accomplished in an instant what he could not do himself. It was, he said, a “weird feeling” (“oh man here it is moment” Altman said referring to the launch of GPT-5), a fleeting but powerful moment of displacement and wonder, one that echoes the experiences of countless others who have, in recent months, tested the boundaries of their own understanding in conversation with these new artificial interlocutors [^2] Specifically the Von interviewer after Altman story did this observation: “Yeah I think that’s I think that feeling right there that’s the feeling a lot of people kind of have like what’s going you know when does it happen what’s going to happen”; see https://www.youtube.com/watch?v=aYn8VKW6vXA. [315].
@@
In another recent interview between Andrew Mayne and Sam Altman, the term “moment of AGI” explicitly appears in the conversation [272]. Mayne observes that many users report experiencing such a moment when interacting with advanced agentic systems, such as Operator with o3333Operator is an AI agent developed by OpenAI that can autonomously control the user’s computer and browser to perform complex, multimodal tasks. The underlying model, “o3”, is OpenAI’s next-generation reasoning engine, designed to integrate external tools and operate on text, images, and live data in real time [287, 213]., describing the impression of watching an AI “use a computer pretty well” [^4] See https://www.youtube.com/watch?v=DB9mjd-65gw%7D%7D. For Mayne, his own “moment of AGI” occurred with Deep Research, as he witnessed the system autonomously gather and synthesize information “better than I read before”, closely resembling a genuinely agentic investigative process. Altman, while noting that the effect was less pronounced for him personally, acknowledges that such moments are becoming widespread as AI capabilities evolve. These firsthand accounts help illustrate how users encounter a threshold of surprise and cognitive resonance while interacting with such smart AI systems [^5] We’re aware of the commercial logic behind this type of interview. However, this doesn’t negate the possibility of intense surprise during a specific interaction with a generative AI system..
@@
What is it that lies at the heart of such experiences? The first encounter with a generative language model is (in the near future equipped with strong agentic capabilities), for many, a moment charged with surprise—a sensation that one’s own thoughts, intentions, or even doubts have been grasped, anticipated, or resolved by a machine. This “wow effect”, as it has come to be called, is not merely a function of technical proficiency, but of a much deeper alignment between linguistic performance and human cognitive expectation. The LLM does not simply return information; it emulates a form of sense-making that appears, at least in the moment, to rival or even surpass that of its human partner [183, 203]—Fig 1 is a pictorial illustration of the “wow effect”.
@@
It is worth noting that the shock of this encounter is amplified by the “cognitive ergonomics” of language. Unlike earlier machines—which demanded users adapt to their logic and limitations—these models adapt themselves, flexibly and contextually, to human expression. This interaction, deeply embedded in natural language, fosters a sense of intimacy and mutual understanding, an illusion of shared meaning that far exceeds anything previously produced by mechanical or classical algorithmic processes [319, 325]. As recent studies have shown, users frequently report feelings of connection, empathy, or even trust in their first interactions with LLMs, especially when the system provides responses that feel both novel and deeply relevant [197, 325].
@@
In attempting to explain this effect, one must take into account the technical details of neural architectures or the mathematical intricacies of transformer models. But there is, at play, an irreducible complexity—a layering of signs [210, 291], meanings, and expectations—rooted in both the architecture of the system and the architecture of human cognition. The user’s sense of surprise is not just a response to statistical fluency or syntactic correctness; it arises from a felt encounter with what seems like a mind at work, a co-constructor of meaning whose origins and intentions remain, in the end, hidden behind the black box of contemporary AI [258, 203]. This leads to the era of “opaque machines”.
@@
The opacity, or “explanatory gap”, is itself not new in the philosophy of mind. As Chalmers argues, the explanatory gap refers to the persistent difficulty in bridging the divide between the observable phenomena—such as linguistic sense or behavioral competence—and the underlying physical or computational mechanisms that give rise to them [187]. Yet, in the case of LLMs, it is compounded by the scale and hierarchy of the models, the astronomical size of their representational space due to the size of the context windows and their ability to generate not just plausible, but sometimes astonishingly creative, responses [182, 329, 178]. The dialogue with an LLM, especially at first contact (or for a user with no technical experience), can take on the character of a magical interaction, reminiscent of the anthropological notion of “magic” explored in the past by Mauss, Lévy-Bruhl, and other classical theorists [271, 255]. The user, unable to reconstruct the causal chains underlying the model’s output, finds herself in a position akin to that of a spectator at a prestidigitation show: she knows, rationally, that there is a “trick”, a mechanism, but is nonetheless enchanted by the effect [203]. The spectator, who today has at least an average level of education, rationally “knows” that the predictor has a suite of magic tricks, yet in most cases remains surprised and in a state of doubt. Moreover, just as audiences willingly suspend disbelief and experience authentic emotional reactions when engaging with fictional characters in literature or film, despite being fully aware of their constructed nature, users of generative AI systems may find themselves empathizing with and attributing interiority to these artificial agents. Empirical studies have shown that such “narrative empathy” and the capacity for simulation are robust features of human cognition [265, 228, 286]. The boundaries between illusion and reality, between mind and mechanism, become fluid and negotiable.
@@
On the other hand, Georg Simmel, in his analysis of modernity at the beginning of XX century, remarked on the ways in which knowledge is increasingly compartmentalized, encapsulated within technical systems whose workings become inaccessible to the layperson [308]. In the age of generative AI, this dynamic is raised to new heights; the black box is not just an engineering metaphor, but a lived phenomenological reality, one that reshapes our experience of agency, knowledge and self. As Altman’s anecdote illustrates, the user may experience both fascination and a kind of existential unease—wonder at the machine’s apparent intelligence and uncertainty about one’s own place in the emerging ecology of minds [176].
@@
It is in response to this new condition that we introduce the concept of “Noosemia”, as a distinct specification of Dennett’s “intentional stance” [207], that is adopting a strategy of interpreting an entity’s behavior by ascribing beliefs and desires to it. The term is designed to capture the distinctive mode of mind attribution and sense-making that arises in human interaction with generative AI. Noosemia is not reducible to classical anthropomorphism or animism. Rather, it is a projection of agency, intentionality, and even interiority, triggered by the linguistic and semiotic performances of the machine and shaped by the interplay of epistemic opacity, creative resonance and dialogic surprise [207, 319, 203]. The phenomenon is not merely a cognitive illusion, but the expression of a new form of “psychological glue”—a spiraling dynamic of fascination, engagement, and dependency, fostered by the machine’s capacity to produce meaning in ways that align with, or exceed, human expectation [319, 181]. This dynamic of engagement is deeply intertwined with what semioticians and philosophers have termed the “surplus of meaning”—as for example U. Eco or U. Galimberti [210, 221], wherein each interaction with the machine generates more interpretive possibilities than can be fully anticipated or exhausted [211, 300]. The encounter with generative AI thus increasingly takes on the structure of a symbolic relationship, in the literary and philosophical sense, where the production and negotiation of meaning are open-ended, dialogic and perpetually oriented toward horizons that transcend any single interpretation [300, 210].
@@
A useful historical precedent for understanding Noosemia is the so-called “Eliza Effect”, a term coined (presumably in the early 1990s) by Douglas Hofstadter in the preface “The Ineradicable Eliza Effect and Its Dangers” of his book Fluid Concepts and Creative Analogies [232]. The Eliza Effect refers to the persistent human tendency to attribute genuine understanding and intentionality to computational systems whose actual capabilities are far more limited than they appear. The original ELIZA program, developed by Joseph Weizenbaum in the mid-1960s [331], was a primitive symbolic system designed to emulate a Rogerian psychotherapist through simple pattern-matching of linguistic input. Despite its simplicity, ELIZA famously elicited strong attributions of understanding and empathy from many users, demonstrating how easily humans project meaning onto machine output. Building on this tradition, the phenomenon of Noosemia can be seen as a contemporary evolution of the “Eliza Effect”, but with fundamental differences. Whereas the “Eliza Effect” was rooted in the response to narrowly scripted, symbolic dialogue systems, Noosemia emerges in interaction with generative AI models—based on connectionist, neuro-symbolic, or agentic architectures—that can emulate a far broader spectrum of expressive and cognitive behaviors, mimicking—for some extent—human creativity. Modern systems are no longer confined to simple text-based exchanges, but can engage in multimodal, context-sensitive, and seemingly open-ended dialogue, often displaying surprising semantic coherence and adaptive reasoning. This expansion of expressive capacity fundamentally transforms the projection dynamic. Hence, Noosemia is not limited to superficial linguistic mimicry, but involves the attribution of agency, intentionality, and even a sense of interiority to systems whose internal logic remains greatly opaque.
@@
However, Hofstadter’s caution remains instructive [^6] Hofstadter on the long run “Eliza Effect” stated [232]: “[…] like a resilient virus capable of continuous mutation, the Eliza effect seems to be resurfacing in the field of AI with ever new disguises and in ever more subtle forms”.. He observed that, despite public enthusiasm, there is a profound gap between genuine conceptual understanding and mere symbolic manipulation. As he famously concluded, “We still have a long way to go, in the scientific study of how the human mind works, before we can build a computer capable of producing even a good joke, let alone a novel” [232]. Yet, as generative models continue to advance, it is increasingly plausible (yet not certain) that, within a few years, LLMs may achieve the capacity to write coherent and compelling novels—reshaping both the landscape of artificial intelligence and our own criteria for meaning and creativity.
@@
As these systems continue to advance [183, 311, 338], the experiences of surprise, fascination, and interdependence that mark noosemic interaction will likely become more frequent and intense, and also increasingly consequential for individuals and society at large. Incidentally, LLMs, according to a study by METR (Model Evaluation & Threat Research), are doubling their capacity to complete complex tasks every seven months [337]. Overcoming limitations is leading to a significant procurement of hardware and computing resources as well as energy. Billion-dollar investments are increasing and do not appear to be reaching a plateau as of today (2025) [269, 327, 268]. We believe this hype is indirectly due in part to a noosemic effect, which has made technologists and investors enthusiastic and confident (or perhaps overconfident) in the AI race and states in redefining geopolitical balances [220, 235]. By articulating and analyzing this phenomenon, we aim to foster a more critical, reflective and imaginative engagement with the emerging “ecology of minds” [176] brought about by generative AI. It is worth noting that the overall purpose of this study is to focus on a specific phenomenon that will be described and analyzed phenomenologically and placed within a pre-existing theoretical framework. However, the study does not operationalize noosemy. The search for testable hypotheses or methodological frameworks for measuring the phenomenon is beyond the scope of this study, which is exploratory in nature.
@@
The first episode concerns the first conversation ever with ChatGPT in 2022; at that time, the available version was the relatively limited GPT-3.5 [244, 182]. The author777i.e., Enrico De Santis. engaged the system by asking for a technical explanation of its own workings, enabling it to converse and explain technical details. For the first time, the author witnessed a machine—though still little more than a “stochastic parrot” [177]—responding in natural language with a command of dialogue and an emulation of comprehension previously unseen in deterministic, passive interfaces. This marked a genuine “quantum leap” compared to previous machine interactions [^8] Even philosopher David Chalmers, a central figure in debates on consciousness and artificial intelligence, admits in a recent interview [226] to being “impressed” by the progress of LLMs such as GPT-3. He notes that we have moved from pure speculation to daily interaction with systems that make once science-fictional scenarios—such as passing the Turing test—plausible realities. See https://shorturl.at/NZJ0k..
A second experience worth sharing involved a test of the first multimodal GPT-4V [183, 306] model, capable of analyzing images natively. The author demonstrated its abilities to a friend working in a café, who was entirely unaware of the recent advances. The author pointed his phone at the inside of a fridge behind the bar—focusing on the bottles inside, visible through the transparent glass door—to see whether GPT-4V could recognize anything. The model not only identified the objects and their brands but also inferred that the setting was a bar, explaining that it “noticed the classic bar counter reflected in the glass”—a detail that was, indeed, present. I observed the unmistakable “spark in my friend’s eyes”, a sudden reaction of surprise and disbelief, that is the very essence of the noosemic experience.
@@
A final episode worth mentioning occurred when the author demonstrated that ChatGPT (the old version 3.5) was capable of “creating” a “metasemantic” poem in the style of the Italian poet Fosco Maraini, renowned for his poem “Il lonfo”[266]. Maraini’s metasemantics seeks to go beyond conventional meaning. The words in “Il lonfo” appear incomprehensible when considered individually, yet, when combined, they evoke a sense of possible meaning in the reader or listener, as their sounds intuitively suggest images or associations. Remarkably, the language model composed a credible piece, displaying compositional abilities at the sub-word level and an implicit understanding of the phonetic qualities of words. An elderly professor of literature who witnessed the experiment was astonished upon discovering that the composition did not exist anywhere else, but was in fact a genuine, statistically generated metasemantic creation.